DIFFERENT POLARIZED COMPONENTS SHAPE OF RAMAN 1343 CM-1 BAND FOR NITRO-BENZENE AND ITS SOLUTIONS

F. H. TUKHVATULLIN, \underline{A. JUMABOEV}, U.N. TASHKENBAEV, S.A. OSMANOV, Z. MAMA-TOV, H. A. HUSHVAKTOV, Samarkand State University, 703004 Samarkand, Uzbekistan.

In Raman spectra for 1343 cm-1 band of liquid nitro-benzene the maxima position of perpendicular and parallel polarized components are differed on =1,8 cm-1. At dilution of nitro-benzene in hexane (within the limits of solubility), chloroform and nitro-methane the frequency difference is decreased and is reduced to zero at the strong dilution. Peculiar kind of dependence the as a function of concentration for mixtures nitro-benzene with nitro-methane is associated with large dipole moment of solvent molecules.